# Yongyang Cai  Khyati Malik   Hyeseon Shin

### v2

@@ 6 Results
In this section, we discuss our simulation results. We first present the results of the noncooperative model under the baseline emission cap scenario for 2020 to 2100. Next, we compare the economic outcomes of different models: noncooperation versus cooperation, with or without the emission caps, and with or without the global ETS. Lastly, we present simulation results with the different emission cap scenarios assuming the net zero target year to be 2050, 2070, and 2090, respectively, for all countries.
6.1 Noncooperation with ETS under the Baseline Emission Caps
@@
Figure 2 displays simulation results at the global level, from the noncooperative model with the ETS under the baseline emission cap scenario. Until 2100, we can delineate three periods according to the global emission trading patterns. Prior to 2030, global emissions do not reach the global emission cap and therefore the permit price remains zero, implying an excess supply of emission permits in the first decade (Figure 2 top-left and bottom-left). This result is not surprising, considering that an oversupply of emission permits has been observed in the EU ETS (Fuss et al., 2018), resulting in zero or very low permit prices. During this period, the volume of global emission abatement and its associated abatement cost remain at low levels (Figure 2 top-left and bottom-right). This result implies that the global emission cap should be set at a level such that there is no over-supply of emission permits, so that permit prices are strictly positive under noncooperation and the ETS.
@@
Global emissions are constrained by the global emission caps starting from 2030. To comply with the monotonically decreasing emission caps, the regions undertake additional abatement efforts and/or purchase emission permits. This results in a substantial increase in the volumes of both emission abatement and traded permits, along with their associated costs. The decrease in global emissions is mainly achieved by the concomitant increase in global abatement, which peaks by around 2065 (Figure 2 top-left). The global abatement cost also increases steeply to reach its maximum value of \$5.18 trillion by 2070, and then decreases as the world reaches net zero emissions by 2070. The emission permit trade, which rises from 2020, gradually decreases to zero by 2070 (Figure 2 top-left). As the emission cap becomes tighter over the years, the permit price increases to \$811 per ton of carbon in 2050 and \$2,254 in 2069 (Figure 2 bottom-left). The kink in the permit price path in 2050 is a result of some regions achieving net zero emissions. As shown in Figure A.2, under the baseline emission cap scenario, the net zero target year is 2050s for Japan (2050) and EU (2053), 2060 for Russia and China, and 2070 for the rest of regions. Essentially, binding emission caps lead to a rise in the overall emission abatement, while the steep increase in the permit price limits the trading of emission permits.
@@
Figure 2: Simulation results at the global scale for the noncooperative model with emission trading under the baseline emission cap scenario.
@@
Post-2070 is the period when global emissions are net zero. For this period, emission permits are no longer traded and all emissions are abated in each region. [^10] In the bottom-left panel of Figure 2, we plot the emission permit price only when the traded volume is positive. By the end of this century, the atmospheric temperature is projected to reach 1.7 degrees Celsius above the pre-industrial level, which is driven by net positive emissions before 2070 (Figure 2 top-right), restricted by the emission caps. [^11] If there is no emission cap, then the temperature anomaly will be much higher. This can be seen in a later discussion with alternative emission caps. Lastly, the global climate damage costs rise almost linearly over the entire period (Figure 2 bottom-right). Overall, our noncooperative model simulation predicts that the global emission caps set by the Glasgow Pact are not restrictive enough to achieve the global target of limiting the temperature rise to 1.5 degrees Celsius.
@@
It is instructive to examine the trading patterns of different regions in the emission permit market over the years to identify the permit buyers and sellers. Figure 3 displays the volumes of traded emission permits for each region over time, with a positive value denoting permit purchase and a negative value indicating permit sales. In the first decade, although there is an excess permit supply at the global level, the emission cap constraint is effective for some regions, such as the United States and Latin America, as these regions become permit buyers at this early stage. After 2029, when the global emission cap constraint becomes binding, the group of permit buyers expands to the US, EU, India, Africa, and OHI regions, with the US, EU, and India being the largest permit buyers. The group of permit sellers consists of Russia, China, Eurasia, OthAs, and Latin America, with China being the largest permit provider. Japan is expected to be involved in a relatively small volume of permit trading.
@@
Figure 3: Simulation results of regional emission permit purchase for the noncooperative model with emission trading, under the baseline emission cap scenario.
Based on our simulation results, we now discuss the optimal regional carbon tax under the noncooperative model with the ETS and the baseline emission cap scenario. Figure 4 shows the regional MAC along with the market equilibrium price of emission permits. We find that the MAC of every region increases drastically and remains strictly greater than the permit price until 2050, and the net emissions are strictly positive for all regions during this time.  [^12] For regional optimal net emissions, see Figure 3 in Appendix A.8.1. As discussed in Section 4.2, $\mathrm{MAC}(E_{i,t}^{*})>m_{t}$ is equivalent to $E_{i,t}^{*}<E_{i,t}^{m}$, implying that when there is no carbon tax, the optimal net emissions for the social planner do not align with the optimal net emissions for the representative firm in each region. Therefore, the optimal carbon tax can be determined by $\tau_{i,t}=\mathrm{MAC}(E_{i,t}^{*})-m_{t}$, so that the firm’s net emissions are equal to the social planner’s optimal net emissions $E_{i,t}^{*}$, for region $i$ at time $t$ before 2050.
@@
Figure 4: Simulation results of regional MAC for the noncooperative model with emission trading under the baseline emission cap scenario.
@@
In Figure 4, the MAC of each region gradually decreases below the permit price after its net zero emissions (before trading of permits) is achieved. Russia is the first region to achieve net zero emissions (before trading of permits) in 2051 (as shown in Figure A.6 in Appendix A.8.1). But Russia still sells permits afterwards, as shown in Figure 3, because its after-trade net zero emission target year is 2060 under the baseline emission cap scenario. After 2051, Russia’s MAC declines below the permit price, because its emission control rate has reached its upper bound. Similar trends are followed by other regions as they attain zero net emissions (before trading of permits) in the 2050s (China, Latin America, and the OHI) and the 2060s (the US, the EU, Japan, Eurasia, India, and MidEast). Africa, and non-OECD Asia are the last group of regions to achieve net zero emissions (before trading of permits) by 2070, after which emission permits are no longer traded, and the MACs of all regions decline slowly over time. [^13] When emission control rates are one, the MACs are $1,000b_{2,i}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))$, and they are nearly constant when $t$ is large. Recall that, when the regional net zero emissions (before trading of permits) are achieved, the optimal carbon tax should be larger than the gap between MAC and permit price, i.e., $\tau_{i,t}\geq\mathrm{MAC}(E_{i,t}^{*})-m_{t}$. Note that if a region’s after-permit-trade zero emission cap constraint is binding, then even when its regional MAC is smaller than the permit price, the region will not be able to sell emission permits, otherwise it will violate the constraint.
@@
Figure 5: Simulation results of regional SCC for the noncooperative model with emission trading under the baseline emission cap scenario.
@@
We further confirm the relationship between the optimal regional carbon tax and the regional SCC under noncooperation with the ETS. Figure 5 shows that when the net emissions of a region are strictly positive, the regional SCC is exactly equal to the optimal carbon tax, which is equal to the gap between the regional MAC and the permit price in Figure 4. For example, in 2050, the MAC for the US is \$971 per ton of carbon, the permit price is \$811, and their difference is exactly equal to the regional SCC of the US, \$160. Figure 5 also shows that when the regional net emissions is zero, the regional SCC is larger than the gap, which could be negative. Among all the regions, the US has the highest SCC, starting at \$106 per ton of carbon in 2020 and steadily increasing up to \$213 per ton of carbon by the end of this century. The US is followed by the EU (from \$42 to \$139 per ton of carbon) and China (from \$39 to \$117 per ton of carbon) in this century. The regions with the lowest SCC in 2100 are Japan (\$29 per ton of carbon), Russia (\$18 per ton of carbon), Eurasia (\$12 per ton of carbon), and MidEast (\$33 per ton of carbon). Our results show that there is considerable heterogeneity in the SCC across regions and suggest that a significant amount of regional carbon tax implementation is required, particularly for the US, the EU, and China, under the global ETS with the baseline emission cap scenario.
6.2 Comparative Analysis of Different Cases
We now modify our baseline model with various cases to understand the implications of cooperation among regions and the ETS. We compare outcomes between the noncooperative and the cooperative models, and for each model, we examine the impact of the ETS under the baseline emission caps. Figure 6 displays the global emission and temperature time profiles for all the cases. Note that for the cooperative model with the ETS, we use the equilibrium solution of permit prices obtained under the noncooperative model with the ETS, so we can compare the difference in the solutions between cooperation and noncooperation, holding permit prices fixed across scenarios.
Figure 6: Simulation results of different cases under cooperation or noncooperation.
We observe that under cooperation, with the objective of maximizing the sum of discounted utility of all the regions, emissions and temperature rises are lower than their counterparts under noncooperation. Specifically, cooperation results in approximately 0.17 degrees Celsius lower temperature with the ETS and 0.13 degrees Celsius lower temperature without the ETS. Our findings underscore that global cooperation is still helpful in mitigating climate change even when every region has the baseline emission caps. Moreover, the ETS has little impact on the temperature path, by reducing the temperature anomaly in 2010 with only 0.02 degrees Celsius, as the emission caps exist no matter whether the ETS exists or not.
When comparing the cases with and without the ETS, [^14] Note that the emission caps are still there, no matter whether the ETS is implemented or not. we find that the ETS may play different roles under noncooperation and cooperation. In the noncooperative model, implementation of the ETS results in slightly higher emissions and temperature rises compared to the case without the ETS. This is because with the ETS, regions with binding regional emission cap constraints now have the option to purchase permits from regions with less restrictive emission caps, fully exploiting the total amount of permits allowed at the global level. This happens until 2050. In the cooperative model, on the other hand, the impact of the ETS on global emissions and temperature changes over time. With the ETS, lower levels of global emissions are observed after 2030. This evolving pattern suggests that the ETS serves as a channel to redistribute emissions, such that the sum of discounted utility of all regions can be maximized over time.
We also compare the regional MAC and SCC across the different cases. Figure 7 presents the regional MAC and SCC under both the cooperative and the noncooperative models with or without the ETS, using the US as an illustrative example. [^15] For comparisons of other regions, refer to Appendices A.8.4 and A.8.3. Under cooperation, the MAC of the US (Figure 7 top-left) is very large in the initial years. This is because the US, as a developed region, needs to mitigate a significant amount of emissions so that developing regions can mitigate less (as shown in Figure A.9 in Appendix A.8.3), thus improving global social welfare. After the 2030s, however, the US net emissions become zero in the cooperative case [^16] For regional net emissions under cooperation and noncooperation, see Appendix A.8.2. and cannot be further mitigated, so its MAC starts to decline over time. The effect of the ETS under cooperation depends on the magnitude of the MAC (see Figure A.9 in Appendix A.8.3): if a region is developed but its MAC is high, then in the ETS, the region would abate less as it can purchase permits (e.g., the US); if a region is developed but its MAC is not too high, then in the ETS, it would abate more (and buy permits at the same time) to improve global social welfare (e.g., the EU); if a region is developing but has small MAC and loose cap constraints, then in the ETS, it would abate more so the region can sell permits with a higher return (e.g., LatAm). Moreover, the ETS has little impact on the net emissions and the MAC of a developed region as shown in Figures A.7 and A.9 (in Appendices A.8.2 and A.8.3) respectively, because a developed region will abate significantly no matter whether the ETS exists or not under cooperation. That is, the ETS plays an important role only in developing regions under cooperation.
Figure 7: Simulation results of different cases for the US in the cooperative or noncooperative model.
The top-right panel of Figure 7 shows that the cooperative SCC of the US is sustained at a high level throughout all periods. This is consistent with the high MAC of the US. When the US regional net emissions are not binding at zero before 2030 , the regional SCC is equivalent to the optimal regional carbon tax, which is also equal to the regional MAC, as the exogenous permit price (from the noncooperative model’s solution) is zero before 2030. Moreover, from our definition of cooperative SCC, we can see that a developed region will have a high SCC, as the marginal global welfare over capital, $\partial V_{t}^{\mathrm{Coop}}/\partial K_{i,t}$, for the developed region is relatively small, while the marginal global welfare over emissions, $\partial V_{t}^{\mathrm{Coop}}/\partial\mathcal{E}_{t}$, is the same across regions. This also implies that a developed region will contribute more to mitigating emissions under cooperation. This pattern is consistent across regions: Figure A.11 in Appendix A.8.4 shows that the US, EU, Japan and OHI have a high cooperative SCC, while the developing regions like India and Africa have a small cooperative SCC.
Under noncooperation, the MAC of the US is substantially lower in the case with the ETS compared to the case without the ETS (Figure 7, bottom-left panel). Such discrepancy persists until 2058, the period during which the US is a net permit buyer (shown in Figure 3), reflecting the fact that implementation of the ETS offers the US a more cost-effective means of keeping its net emissions below its cap. Consequently, the US conducts less emission abatement. For all other regions, we find that the MAC is consistently lower (higher) with the ETS compared to the case without the ETS when a region is a buyer (seller) across all periods with nonzero permit prices from 2030 (see Figure A.10 in Appendix A.8.3).
The bottom-right panel of Figure 7 shows that the noncooperative SCC of the US is much smaller than its cooperative SCC, which is also consistent with the smaller MAC of the US under noncooperation. Moreover, every region has a much smaller SCC under noncooperation than under cooperation, as shown in Figures A.11 and A.12 in Appendix A.8.4. This result implies that, in the dynamic Nash equilibrium under noncooperation, no region is willing to impose a high carbon tax thereby allowing other regions to free ride on lower emissions and climate damages. Therefore, in the noncooperative model, emissions and temperature rises are much larger than in the cooperative model, as shown in Figure 6.
Now we confirm the relationship between the optimal regional carbon tax and the regional SCC when there is no ETS. For the cooperative case without the ETS, the US regional net emissions are not binding at their caps, but are binding at zero after 2033 (see Figure A.7), so the regional SCC is larger than the regional MAC after 2033 (see the two top panels of Figure 7). For the noncooperative case without the ETS, the US regional net emissions are binding at their caps after 2022 (see Figures A.2 and A.8), so the regional SCC is smaller than the regional MAC after 2022 (see the two bottom panels of Figure 7). Moreover, under no ETS, if the regional net emissions are not binding at their caps or zero, then the regional SCC is equal to the regional MAC (see, e.g., Eurasia under cooperation until 2060 in Figures A.7, A.9, A.11; and Eurasia under noncooperation until 2049 in Figures A.8, A.10, and A.12), which is equal to the optimal regional carbon tax. This relationship holds for every region as shown in Figures A.7-A.12.
The two right panels of Figure 7 collectively demonstrate that the ETS has relatively little impact on the regional SCC under cooperation or noncooperation. This pattern holds for every region as shown in Figures A.11 and A.12. This happens under noncooperation because the regional noncooperative SCC is the present value of of future climate damages in the region resulting from an additional unit of global emissions released in the current period, and the ETS has little impact on the marginal damages and our endogenous social discount rates. Under cooperation, since the ETS has little impact to developed regions and the temperature path, the wealth redistribution among developing regions via the ETS has little impact on the marginal global welfare over capital, $\partial V_{t}^{\mathrm{Coop}}/\partial K_{i,t}$,, so the ETS has little impact on the regional cooperative SCC. This shows that the regional SCC is complementary to the ETS, as the emission caps exist no matter whether the ETS exists or not.
6.3 Alternative Net Zero Scenarios

### v3

@@ 6 Results
In this section, we discuss our simulation results. We first present the results of the noncooperative model under the baseline emission cap scenario for the period from 2020 to 2100. Next, we compare the economic outcomes across different models, examining the role of the emission caps and the implementation of a global ETS. Lastly, we present simulation results with the different emission cap scenarios assuming the net zero target year to be 2050, 2070, and 2090, respectively, for all 12 regions.
6.1 Key Results under the Baseline Emission Caps
@@
Figure 1 displays simulation results at the global level, from the noncooperative model with the ETS under the baseline emission cap scenario. Until 2100, we can delineate three periods according to the global emission trading patterns. Prior to 2025, global emissions do not reach the global emission caps and therefore the permit price remains zero, implying an excess supply of emission permits in the initial years (Figure 1 top-left and bottom-left), as the emission caps in the initial years are large while they become much smaller over time (Figure A.5). This result is not surprising, considering that an oversupply of emission permits has been observed in the EU ETS (Fuss et al.,, 2018), resulting in zero or very low permit prices. This is also consistent with our theoretical analysis for the regional social planners’ Nash equilirbium in Section 2. During this period, the volume of global emission abatement and its associated abatement cost remain at low levels (Figure 1 top-left and bottom-right). This result implies that the global emission cap should be set at a level such that there is no over-supply of emission permits, so that permit prices are strictly positive under noncooperation and the ETS.
@@
Global emissions are constrained by the global emission caps starting from 2025. To comply with the monotonically decreasing emission caps, the regions undertake additional abatement efforts and/or purchase emission permits. This results in a substantial increase in the volume of emission abatement, along with the abatement cost and permit price. The decrease in global emissions is mainly achieved by the concomitant increase in global abatement, which peaks by around 2060 (Figure 1 top-left). The global abatement cost also increases steeply to reach its maximum value of \$5.45 trillion by 2070, and then decreases as the world reaches net zero emissions by 2070. The emission permit trade gradually decreases to zero by 2070 (Figure 1 top-left). As the emission cap becomes tighter over the years, the permit price increases to \$845 per ton of carbon in 2050 and \$2,059 in 2069 (Figure 1 bottom-left). The kinks in the permit price path in 2050 and 2060 are a result of some regions achieving net zero emissions, as shown in Figure A.5. Essentially, binding emission caps lead to a rise in the overall emission abatement, while the steep increase in the permit price limits the trading of emission permits.
@@
Figure 1: Simulation results at the global scale under the baseline emission cap scenario.
@@
Post-2070 is the period when global emissions are net zero. For this period, emission permits are no longer traded and all emissions are abated in each region. [^16] In the bottom-left panel of Figure 1, we plot the emission permit price only when the traded volume is positive. By the end of this century, the atmospheric temperature is projected to reach 1.7°C above the pre-industrial level, which is driven by net positive emissions before 2070 (Figure 1 top-right), restricted by the emission caps. [^17] If there is no emission cap, then the temperature anomaly will be much higher. This can be seen in a later discussion with alternative emission caps. Lastly, the global climate damage rises almost linearly over the entire period (Figure 1 bottom-right). Overall, our noncooperative model simulation predicts that the global emission caps set by the Paris Agreement and Glasgow Pact are not restrictive enough to achieve the global target of limiting the temperature rise to 1.5°C.
@@
It is instructive to examine the trading patterns of different regions in the emission permit market over the years to identify the permit buyers and sellers. Figure 2 displays the volumes of traded emission permits for each region over time, with a positive value denoting permit purchase and a negative value indicating permit sales. Before 2025, although there is an excess permit supply at the global level, the emission cap constraint is effective for some regions, such as China and Latin America, as these regions become permit buyers at this early stage. After 2025, when the global emission cap constraint becomes binding, the group of permit buyers consists of the US, EU, India, and OHI regions; the group of permit sellers consists of Russia, China, and Eurasia, with China being the largest permit provider after 2036. Japan is expected to be involved in a relatively small volume of permit trading, and the other regions (Africa, MidEast, Latin America, and OthAs) change their status of permit suppliers to buyers or vice versa over time.
@@
Figure 2: Simulation results of regional emission permit purchase under the baseline emission cap scenario.
6.2 SCC, MAC, and the Optimal Carbon Tax
Based on the simulation results under the baseline emission cap scenario, this section examines the relationship between the MAC, regional social cost of carbon (SCC), and the optimal carbon tax under the global ETS regime. The MAC captures the additional cost incurred due to an increase in emission abatement. From our model equation (10), the total abatement cost is $\mathrm{TAC}{}_{i,t}=b_{1,i,t}\mu_{i,t}^{b_{2,i}}Q_{i,t},$ in trillions of USD. Thus, the MAC, in 2020 USD per ton of carbon, is obtained as follows:
$\mathrm{MAC}{}_{i,t}=1,000\left(\frac{\partial\mathrm{TAC}{}_{i,t}}{\partial E_{i,t}^{A}}\right)=1,000\left(\frac{b_{1,i,t}b_{2,i}\mu_{i,t}^{b_{2,i}-1}}{\sigma_{i,t}}\right).$ (29)
With $b_{1,i,t}>0$ and $b_{2,i}>2,$ MAC is strictly increasing on emission control rate $\mu_{i,t}\in[0,1].$ [^18] See Table A.2 in Appendix A.1 for the list of our calibrated abatement cost parameters. Figure 3 displays the regional MAC along with the market equilibrium price of emission permits. We find that, for every region, the MAC increases rapidly and remains strictly greater than the permit price until the emissions hit zero by 2050s or 2060s. [^19] For regional optimal emissions, see Figure 11 in Appendix A.4.1. As discussed in Section 2, $\mathrm{MAC}(E_{i,t}^{*})>m_{t}$ implies that the optimal emissions for the regional social planner do not align with the optimal emissions for the representative firm under no tax in each region. Therefore, the optimal regional carbon tax can be introduced at $\tau_{i,t}=\mathrm{MAC}(E_{i,t}^{*})-m_{t}$, so that the firm’s emissions are equal to the regional social planner’s optimal emissions $E_{i,t}^{*}$, for region $i$ at time $t$ before 2050 or 2070. Russia is an exception, where the MAC falls below the permit price starting in 2025, even though it has nonzero emissions until 2050, because global warming leads to increase in productivity rather than causing climate damages in Russia. This result is strongly influenced by the climate damage parameters, calibrated using projections from Burke et al., (2018).
@@ 6.2 SCC, MAC, and the Optimal Carbon Tax
Figure 3: Simulation results of regional MAC under the baseline emission cap scenario.
@@
In Figure 3, the MAC of each region gradually decreases below the permit price after its net zero emissions is achieved (i.e., $E_{i,t}^{*}=0$). Russia is the first region to achieve net zero emissions in 2050 (as shown in Figure A.7 in Appendix A.4.1). However, Russia continues to sell permits afterwards, as shown in Figure 2, because its net zero emission target year is 2060, under the baseline emission cap scenario. After 2050, Russia’s MAC begins to decline slightly and stabilizes over time, as its emission control rate reaches its upper limit. [^20] When emission control rates are one, the MACs are $1,000b_{2,i}(b_{1,i}+b_{3,i}\exp(-b_{4,i}t))$, and they are nearly constant when $t$ is large. Similar trends are followed by other regions as they attain net zero emissions in the 2050s ( China, the US, the EU, the OHI, MidEast, Eurasia, and Latin America) and the 2060s (Japan and India). Africa and OthAs are the last group of regions to achieve net zero emissions (before trading of permits) by 2070, after which emission permits are no longer traded, and the MACs of all regions decline slowly over time. Recall that, when the regional net zero emissions (before trading of permits) are achieved (i.e., $E_{i,t}^{*}=0$), the regional optimal carbon tax should be larger than the gap between the regional MAC and permit price, i.e., $\tau_{i,t}\geq\mathrm{MAC}(E_{i,t}^{*})-m_{t}$. Note that if a region’s after-permit-trade zero emission cap constraint is binding (i.e., the inequality (12) is binding with $\overline{E}_{i,t}=0$), then even when its regional MAC is smaller than the permit price, the region will not be able to sell emission permits, otherwise it will violate the constraint.
Given the simulation results for the optimal carbon tax, we now turn to discuss the relationship between the optimal carbon tax and the regional social cost of carbon (SCC). The SCC, a central concept in the climate change literature, is widely used to quantify the present value of climate damages induced by an additional unit of carbon emissions. While the SCC is often calculated in a global social planner’s problem (e.g., the DICE model), we consider the SCC for each region. Similar to van der Ploeg and de Zeeuw, (2016) and Cai et al., (2023), we define the noncooperative SCC of a region as the marginal rate of substitution between global emissions and regional capital as follows:
$\mathrm{SCC}_{i,t}=\frac{-1,000(\partial V_{i,t}/\partial\mathcal{E}_{t})}{\partial V_{i,t}/\partial K_{i,t}},$ (30)
where $V_{i,t}$ is the value function of the noncooperative model at time $t$ for region $i$, depending on the state variables $\{K_{i,t},\mathcal{E}_{t}:\ i\in\mathcal{I}\}$; that is,
$V_{i,t}(K_{1,t},...,K_{12,t},\mathcal{E}_{t})=\max_{c_{i,s},E_{i,s}^{P},\mu_{i,s}}\sum_{s=t}^{\infty}\beta^{s-t}u(c_{i,s})L_{i,s},$
for each region $i$ under the open loop Nash equilibrium. Since our cumulative global emissions are measured in gigatonnes of carbon (GtC) and capital is measured in trillions of USD, our SCC is measured in 2020 USD per ton of carbon. [^21] To compute the regional SCC in the noncooperative model, it is equivalent to replace the numerator in equation (30) with the shadow price of the transition equation of cumulative global emissions (14) at time $t$, and replace the denominator with the shadow price of the regional capital transition equation (20) for each region.
Our concept of the regional SCC differs from the definition in Nordhaus, (2017) and Ricke et al., (2018), where the regional SCC is defined as the present value of future climate damages in a region resulting from an additional unit of global emissions released in the current period, but their social discount rates are exogenous. We define the regional noncooperative SCC as the marginal rate of substitution between cumulative global emissions and regional capital under the Nash equilibrium. With our definition, the regional noncooperative SCC is equal to the optimal regional carbon tax under noncooperation when the regional emissions are not binding at zero (i.e., the regional emission control rate is less than one), or is larger than the optimal regional carbon tax when the regional emissions are binding at zero. This pattern is consistent with the relation between the global SCC and the optimal global carbon tax as shown in Cai et al., (2017) and Cai and Lontzek, (2019).  [^22] We also verified numerically that our regional noncooperative SCC is equal to the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, with our social discount rates $r_{i,t+1}$ defined endogenously with the following formula: $r_{i,t+1}=\frac{u^{\prime}(c_{i,t})}{\beta u^{\prime}(c_{i,t+1})}-1,$ where $c_{i,t}$ are the optimal per-capita consumptions.
@@
Figure 4: Simulation results of regional SCC under the baseline emission cap scenario.
@@
The simulation results confirm the relationship between the optimal regional carbon tax and the regional SCC under noncooperation with the ETS. Figure 4 shows that when the emissions of a region are strictly positive, the regional SCC is exactly equal to the optimal carbon tax, which is equal to the gap between the regional MAC and the permit price in Figure 3. For example, in 2050, the MAC for the US is \$1,081 per ton of carbon, the permit price is \$845, and their difference is exactly equal to the regional SCC of the US, \$236. Figure 4 also shows that when the regional emissions are zero, the regional SCC is larger than the gap, which could be negative. Among all the regions, the US has the highest SCC in the near term, at around \$198 per ton of carbon in 2030. However, the SCCs of Africa and non-OECD Asia are expected to experience substantial increases, reaching \$682 for Africa and \$610 for non-OECD Asia by the end of the century. Russia, with the lowest SCC, experiences a steady negative SCC, reaching (-\$88) per ton of carbon by the end of the century, indicating that it benefits from global warming throughout the entire period. Our results show considerable heterogeneity in the SCC across regions and suggest that significant regional carbon tax implementation is required, not only for developed countries such as the US, but also for developing countries like Africa, India, and non-OECD Asia under the global ETS with the baseline emission cap scenario.
6.3 Effects of ETS Implementation
Next, we examine the economic and climate implications of the ETS implementation by comparing the global economy with and without the ETS regime, while maintaining the baseline emission caps in both economies. The top two panels of Figure 5 show that the ETS implementation results in slightly higher emissions and temperature increases compared to the case without the ETS. This occurs because, under the ETS, regions with binding regional emission cap constraints now have the option to purchase permits from regions with less restrictive emission caps, fully exploiting the total amount of permits allowed at the global level, but the buyers do not have this option under no ETS. That is, the global emissions could be less than the global emission caps under no ETS, while they are equal under the ETS. This pattern persists until 2043. In the bottom two panels of Figure 5, we compare the MAC and the SCC using the US as an illustrative example. The MAC of the US economy is higher without the ETS until 2057, reflecting that the US becomes a permit buyer under the ETS regime as MAC increases with increasing emission abatement. The SCC comparison demonstrates that the ETS has little impact on the regional SCC. This pattern holds for the other regions as well, as shown in Figure A.10. Since the regional noncooperative SCC is the present value of future climate damages in the region resulting from an additional unit of global emissions released in the current period, these results show that the ETS has little impact on the marginal damages and our endogenous social discount rates.
Figure 5: Comparative analysis: effects of the ETS implementation.
What are the welfare implications of implementing the global ETS? To quantify these effects, we compute a compensating variation (CV) per capita associated with the ETS implementation. Specifically, the CV per capita for region $i$ is computed by solving the following equation:
$W_{i,0}(\mathbf{c}_{i}^{1}-CV)=W_{i,0}(\mathbf{c}_{i}^{0}),$ (31)
where $\mathbf{c}_{i}^{1}=(c_{i,0}^{1},\cdots,c_{i,t}^{1},\cdots,c_{i,T}^{1})$ represents the vector of optimal consumption per capita under the ETS implementation, $\mathbf{c}_{i}^{0}=(c_{i,0}^{0},\cdots,c_{i,t}^{0},\cdots,c_{i,T}^{0})$ is the vector of optimal consumption per capita without the ETS implementation, and
$W_{i,0}(\mathbf{c}_{i})=\sum_{t=0}^{T}\beta^{t}u(c_{i,t})L_{i,t}$
is the social welfare associated with the vector of consumption per capita $\mathbf{c}_{i}=(c_{i,0},\cdots,c_{i,t},\cdots,c_{i,T})$ with the terminal time $T=299$. Note that a direct comparison of the noncooperative model with the ETS to the noncooperative model without the ETS does not provide a clear picture of the welfare effects of the ETS under the current baseline emission cap scenario, as the global emissions and resulting climate damages differ across the two economies. That is, the differences between $\mathbf{c}_{i}^{1}$ and $\mathbf{c}_{i}^{0}$ depend on both the emission caps and the ETS implementation. Compared with the economy without the ETS, temperature under the ETS implementation is expected to be higher as shown in Figure 5. Consequently, some regions may experience additional climate damage from the difference of temperatures between two cases, while it is greater than the reduced abatement cost from purchasing permits or the additional profit from selling permits under the ETS. This leads to negative welfare effects of the ETS implementation for some regions. To isolate the impact of the ETS, we adjust the emission caps $\overline{E}_{i,t}$ in the noncooperative model with the ETS to $\overline{E}^{\prime}_{i,t}$, which are the optimal emission levels obtained from the noncooperative model with the emission caps $\overline{E}_{i,t}$ but without the ETS, then we compare economies with and without the ETS under the emission caps $\overline{E}^{\prime}_{i,t}$ as these two economies have the same global emission levels and temperatures.
Table 1 shows the CV per capita in 2020 USD, and its share (%) out of per capita consumption in each region at $t=0$, when the ETS implementation does not impact temperature. The findings indicate that the welfare impacts of the global ETS are strictly positive for all regions, irrespective of whether they are permit sellers or buyers, with considerable heterogeneity in the magnitude of these effects across the regions. For example, in the United States, the CV per capita is \$87.22, which is a relatively small fraction of per capita consumption (0.2%). Among the 12 regions, Russia experiences the largest welfare improvement, with a CV per capita of \$183.53, equivalent to 2.2% of its per capita consumption. In contrast, Africa and non-OECD Asia have the smallest welfare gains from the global ETS. These results highlight that the principle of gains from trade applies to emissions trading as well, driven by efficiency gains achieved through the reallocation of emission abatement efforts across regions.
Table 1: Welfare effects of the ETS implementation.
US EU Japan Russia Eurasia China
CV per capita (\$) 87.22 86.43 35.12 183.53 50.96 24.93
CV (%) 0.201 0.344 0.124 2.225 1.017 0.250
India MidEast Africa LatAm OHI OthAs
CV per capita (\$) 37.09 54.96 13.06 56.93 69.33 8.15
CV (%) 2.374 0.590 0.857 0.865 0.201 0.348
7 Sensitivity Analysis
We perform sensitivity analyses to evaluate alternative emission cap scenarios and to examine the influence of key parameters—climate damages and emission abatement—on our simulation results.
7.1 Alternative Emission Cap Scenarios